In the following electropherogram, for each STR loci identify the heterozygous peak height ratio, the major and minor allele number, and contributor ratio

100 600 200 12 329. 12 27 196.29 1940 32.2 217.85 681 15 12 3934 152.73 4669 722 262.92 279.15 2777 464 309.01 898 13 143.89 28 200.08 10 21.64 271.00 30 207.97 439 123.39 2582 170.95 819 228.40 1143 268.05 392 16 311.30 1433 13 236.28 2151 276.09 1892 19 323.97 950 127.39 74.95 182.92 2127 14 240.29 1601 12 280.18 1573 2,2012 11:24AM, EDT Printed by: gmid
Answer #1

For each STR loci identify the heterozygous peak height ratio,major,minor,contributor ratio is calculated as:

D8S1179- (PHR)-911/4669=0.1951*100=19.51% or 722/911*100=79.2%(approx.)=0.792,

major allele No.-13,15,minor allele No.-12,13 , CR-1633/5580=29.2% or 0.292(approx.)

D21S11-1911/1940*100=98.5%(approx)=0.985 , major-32.2,30,minor=27,28, CR=1120/3851=0.29=29.0%(approx.)

D7S820- 464/480*100=96.6%(approx.)=0.966, major-10,12,minor-8,10, CR-944/3257=28.9%=0.289(approx.)

CSF1PO- 1117/1350*100=82.7%(approx.)=0.827, major=10,12,minor=7,10,CR-2248/2467=91.1%=0.911(approx.)

D3S1358- 2469/2582*100=95.6%=0.956(approx.), major=16,minor=15,CR=0.956(approx.)

TH01- 2127/2996*100=71%=0.7099=0.71(approx.),major-7,9, minor=6,7, CR=3815/5123=74.4=0.744(approx.)

D13S317=1601/2151*100=74.4%=0.744(approx.),major=13,14, minor=11,13, CR=3294/3752=87.8%=0.878(approx.)

D16S539-1573/1892*100=83.1%=0.831(approx.) ,major=11,12, minor=9,11 ,CR=2284/3465=65.9=0.659(approx.)

D2(not visible)-950/1433*100=66.2%=0.662(approx.),major=19,minor=16, CR=0.662(approx.)